OpenStudy (anonymous):

dff wrt x : f=arctan(3y/x)

OpenStudy (asnaseer):

use method as I showed you in your previous question, let w = 3y/x and then differentiate by parts.

OpenStudy (anonymous):

i did that just want to be sure

OpenStudy (asnaseer):

what answer did you get?

OpenStudy (anonymous):

z=arctan(3y/x) w=3y/x then: z=arctan(u) dz/dx=dz/dw∗dw/dx=arctan(u)∗-3yx^-2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@ asnaseer is arctanu= 2arctan[x/1+\[\sqrt{1+x^2}\][3y] ?

OpenStudy (anonymous):

are u dre?

OpenStudy (asnaseer):

sorry I was called away just got back

OpenStudy (asnaseer):

you got dz/dw wrong

OpenStudy (asnaseer):

it should be 1/(1+w^2)

OpenStudy (asnaseer):

\[y=arctan(x)\]\[dy/dx=\frac{1}{1+x^{2}}\]

OpenStudy (anonymous):

Okay. Thanks.